OpenStudy (tester97):

What is the vertex form of the equation? y = –x^2 +12x – 4

OpenStudy (tester97):

@Luigi0210

OpenStudy (luigi0210):

Do you know the vertex formula?

OpenStudy (tester97):

isnt it y=a(x-h)^2 +k?

OpenStudy (luigi0210):

\[\LARGE -\frac{b}{2a}\]

OpenStudy (luigi0210):

\[\LARGE y=ax^2+bx+c\]

OpenStudy (luigi0210):

That will give you the x-coordinate, just plug that back in to find the y-coordinate.

OpenStudy (tester97):

what? do the -b/2a to find the x then plug it into the y=ax^2 +bx+c ?

OpenStudy (luigi0210):

Yup, so find the x-coordinate.

OpenStudy (tester97):

 so it would look like this? and the answer would be 6?